# Structured Reinforcement Learning for Media Streaming at the Wireless Edge

### v1

@@ 1. Introduction
Video streaming has grown in recent years to occupy about 48% of mobile data traffic, with YouTube being the most popular and representing close to 50% of this traffic (Sandvine, 2021). At the same time, increasing softwarization is taking place in both cellular and WiFi domains, allowing fine grained dynamic control of wireless resource usage to support mobile applications via intelligent control (Open Networking Foundation, 2021). Indeed, intelligent control can enable dynamic priority access over the wireless edge (access network) for specific applications that require it, such as a recent demonstration of dynamic network slicing for reliable robot operation over the network (Hewlett Packard Enterprise, 2021).
@@
Figure 1. Feedback loop for reinforcement learning in a media streaming application. The states of the YouTube sessions and channel qualities are communicated to an intelligent controller that determines the service class for each session, accounting for resource constraints. The impact of this decision on the end-user QoE (reward) is communicated back to the controller.
@@ 1.1. Main Results
We first decompose the centralized CMDP problem across all clients into decentralized single-client problems using Lagrangian relaxation techniques, wherein a Lagrange multiplier acts as a price signal to enforce the constraint. As in several other queueing problem contexts, we find that the optimal policy of such a single-client problem has a threshold structure. Specifically, for any given stall count, the decision on whether to assign high priority service to a client only depends on whether or not its video buffer is below a fixed threshold.
We next turn to the important question of learning the optimal policy. Here, the threshold structure of the optimal policy enables us to confine our search to learning the threshold parameter using a single neuron, as opposed to a complex function that needs to be approximated with a large neural network. We also observe that the cost-to-go function for a given stall count is unimodal in the threshold parameter, i.e., the cost increases on both sides of the optimal threshold. This observation allows us to design a three time-scale stochastic approximation algorithm following a policy gradient approach that can converge to the optimal policy. We update the value function in the first time scale, the policy in the second time scale, and the Lagrange multiplier in the third time scale.
We then develop a simulation environment to train optimal policies. We show that the single-client decomposition approach indeed learns at about four times faster than centralized learning while attaining a similar performance. We also develop a heuristic index-variant of our optimal policy that has the advantage of being applicable in spite of a changing number of clients or variability in channel quality.
@@ 1.2. Related Work
There is an extensive amount of work on CMDPs focusing on the problems of control (Altman, 1999, 2002). In the context of media streaming over wireless channels, several works identify structural properties of the optimal policy (ParandehGheibi et al., 2011; Singh and Kumar, 2019; Hsieh and Hou, 2018) that often take a threshold form. While (ParandehGheibi et al., 2011; Singh and Kumar, 2019) consider maximizing QoE for a finite system, and (Singh and Kumar, 2019) utilizes a dual-based decentralization approach, (Hsieh and Hou, 2018) focuses on the heavy traffic limit.
@@
Reinforcement learning approaches to find the optimal policies for CMDP problems are well studied in the recent literature. In the context of tabular setting, recent works focus on characterizing the regret (Efroni et al., 2020; Ding et al., 2021; Liu et al., 2021b; Wei et al., 2022; Bura et al., 2021) and the sample complexity (HasanzadeZonuzy et al., 2021a; Kalagarla et al., 2021; HasanzadeZonuzy et al., 2021b; Vaswani et al., 2022) of learning algorithms. There also a number of works on the policy gradient approaches for CMDP (Achiam et al., 2017; Yang et al., 2019; Zhang et al., 2020; Paternain et al., 2019; Ding et al., 2020b; Liu et al., 2021a; Zhou et al., 2022). However, these works are for the general CMDPs and they do not exploit the structure of the problem to learn a threshold policy with provable global convergence guarantees. The work that is closest to ours is (Roy et al., 2021) which develops a two time scale stochastic approximation algorithm (Borkar, 2009) for an unconstrained MDP, and utilises the structural properties of the optimal policy to attain convergence to the global optimum. Our analytical approach towards exploiting problem structure is motivated by this work. Stochastic approximation algorithm with provable guarantees for constrained reinforcement learning was first proposed in (Borkar, 2005) for the tabular setting. This approach has been extended with function approximation in (Bhatnagar, 2010), but only with local convergence guarantees. In this work, we exploit the structural properties of the underlying problem with the analytical approach of three time scale stochastic approximation (Borkar, 2005) to learn threshold policies with provable global convergence guarantees.
@@
A variety of systems have been proposed to improve the QoE performance of video streaming. In software defined networks (SDN), assigning the video streaming flows to network links according to path selection algorithms is considered in (Jarschel et al., 2013). VQOA (Ramakrishnan et al., 2015) and QFF (Georgopoulos et al., 2013) employ SDN to monitor the traffic and adapt the bandwidth assignment of video flows to achieve better streaming performance in the home network. Reinforcement learning has also been applied to a variety of video streaming applications (Mao et al., 2017; Bhattacharyya et al., 2019), which show significant improvements over existing methods. (Mao et al., 2017) proposes an algorithm Penseive, to train adaptive bit rate (ABR) algorithms to optimize the QoE of the clients. (Bhattacharyya et al., 2019) considers the complementary problem of service prioritization, and shows empirically that off-the-shelf model-based and model-free RL approaches can result in major improvements in QoE. While completely heuristic in its approach, the work motivates us to consider the problem of optimal policy design, while exploiting problem structure. Thus, we provide a simple, structured RL algorithm with provable convergence guarantees, which is empirically able to attain the same performance as more complex deep learning approaches using significantly fewer data samples.
@@ 2. PROBLEM FORMULATION
We consider a discrete time system with $N$ clients. At any time $t$, the media server sends some number of media packets to the AP, which in turns forwards them to the client, which buffers up this content while playing out. The media playout is said to stall when the buffer is empty. Stall events cause end-user dissatisfaction, and we will utilize a standard modal of this disutility in the video streaming context.
@@
We assume that $\mu_{n,1}>\mu_{n,2}$, for each $n$. This means that when the client is assigned to high priority service, its service rate is higher than that of the low priority service. The joint system action space is $\mathcal{A}\triangleq\otimes_{n=1}^{N}\mathcal{A}_{n}$.
Note that we could easily include video bitrate adaptation into the model by simply increasing the action space to multiple dimensions. This causes no changes to the approach. We now describe the dynamics of the system.
@@
State Transition Structure: The system state has two transition processes in its components. The first process corresponds to the media buffer evolution, while the second corresponds to the stall count evolution. We assume that the media playback process of client n, denoted by $B_{n}(t)$ is distributed according to Bernoulli distribution with mean $\beta_{n}<\mu_{n,1},$ i.e., stall-free playout is possible if the client is given high priority service. Also, a stall corresponds to a transition from a non-zero media buffer state to a zero media buffer state.
We assume that the media that is currently being played by the client may be terminated at any time with probability $\alpha>0$ by the end-user, which causes a reset of the client state to $(0,0).$ Hence, the possible state transitions from $s_{n}(t)$ to $s_{n}(t+1)$ are
@@
where $x^{+}=\max(x,0)$, $y_{n}^{\prime}(t)=y_{n}(t)+\mathbbm{1}\{x_{n}(t)=1\}$, and,
${P_{n,1}(a)=1-P_{n,2}(a)-P_{n,3}(a)},P_{n,2}(a)=\mu_{n}(a)(1-\beta_{n}),P_{n,3}(a)=(1-\mu_{n}(a))\beta_{n}$, for $a\in\{1,2\}$, and for every $n$. The state transition probability of the $n$-th client is denoted by $p_{n}(s^{\prime}|s,a)$. The starting state distribution of the $n$-th client is denoted by $\rho_{n}$. Note that we can include client-specific modifications of the transition probabilities based on their individual channels without changing the transition structure.
@@
Note that we choose to use a cost function, rather than a QoE-based reward function for consistency with a generic MDP/RL approach that focuses on cost minimization. It is easy to translate the instantaneous cost to an instantaneous reward by simply multiplying by $-1$ and maximizing. As pointed out above, we can easily include the video bitrate (corresponding to the video resolution) in this cost function but choose not to do so for simplicity. In our experiments, we fix the resolution at 1080p so that all clients obtain a high resolution video stream.
@@
$\min_{\pi\in\Pi}\mathbb{E}^{\pi}[\sum_{n=1}^{N}\sum_{t=0}^{\infty}\gamma^{t}c(s_{n}(t),s_{n}(t+1))]$
@@
(1) $&\min_{\pi\in\Pi}\mathbb{E}^{\pi}[\sum_{n=1}^{N}\sum_{t=0}^{\infty}\gamma^{t}c(s_{n}(t),s_{n}(t+1))]\\ &\text{s.t}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}^{\pi}[\sum_{n=1}^{N}\sum_{t=0}^{\infty}\gamma^{t}g(a_{n}(t))]\leq\bar{K},\end{aligned}$
@@
An immediate consequence of (2) is that there exists a primal-dual pair $(\pi^{*},\lambda^{*})$ achieving strong duality. Denote the individual value and discounted constraint functions of client $n$ under policy $\pi$ by,
(2) $J_{c}^{\pi}(\rho_{n})$ $\triangleq\mathbb{E}_{s_{n}(0)\sim\rho_{n}}[\mathbb{E}^{\pi}[\sum_{t=0}^{\infty}\gamma^{t}c(s_{n}(t),s_{n}(t+1))|s_{n}(0)]],$
(3) $J_{g}^{\pi}(\rho_{n})$ $\triangleq\mathbb{E}_{s_{n}(0)\sim\rho_{n}}[\mathbb{E}^{\pi}[\sum_{t=0}^{\infty}\gamma^{t}g(a_{n}(t))|s_{n}(0)]],$
@@
Given that assumption (2) holds true. Then, $D(\lambda^{*})=\sum_{n=1}^{N}J_{c}^{\pi^{*}}(\rho_{n})$
However, finding the optimal joint policy even for a fixed $\lambda$ in the dual function can be computationally hard because the complexity increases exponentially with the number of users. It can be shown that the centralized problem can be decomposed in to $N$ smaller problems (Singh and Kumar, 2019).
@@
The proof is similar to that of (Singh and Kumar, 2019, Theorem $1$), for completeness, we provide the details in Appendix .2.
Note that Theorem (1) states that it is sufficient to look at decentralized class of policies $\pi=\otimes\pi_{n}$. Henceforth, we refer to the optimal policy as $\pi^{*}=\otimes\pi_{n}^{*}(\lambda^{*})$, and use the shorthand notation for the cumulative objective functions $\bar{J}_{c}^{\pi}\triangleq\sum_{n=1}^{N}J_{c}^{\pi_{n}}(\rho_{n}),\bar{J}_{g}^{\pi}\triangleq\sum_{n=1}^{N}J_{g}^{\pi_{n}}(\rho_{n})$, $\bar{J}_{\lambda}^{\pi}\triangleq\bar{J}_{c}^{\pi}+\lambda(\bar{J}_{g}^{\pi}-\bar{K})$.
3. Structure of the optimal client policy: Threshold and Unimodality
@@ 3.1. Threshold Structure
We will use the following results to prove the above theorem. Since the transition function of all the clients have the same properties (even though they may not be identical), we drop the the client index $n$ for ease of notation, and denote the optimal value function by $J^{*}$ in short hand.
@@
The proofs of Lemma 2 and Lemma 3 are given in Appendix .3. We now present the proof of Theorem 2.
@@ 4. RL for Learning the Optimal Threshold
In this section, we present a primal dual algorithm to solve the constrained markov decision process problem and provide results for its finite time convergence to the global optimal policy. Our implementation of the proposed algorithm follows the methodology of actor-critic algorithms with function approximation. To begin with, we slightly modify our hard-threshold policy to a parameterized soft-threshold policy. For the soft threshold parameterization, the probability of taking action $a$ in state $s$ corresponding to a threshold parameter vector $\theta$, $\pi_{\theta}(s,a)$ is defined as
$\pi_{\theta}(s,a)$ =$=\begin{cases}1-\frac{1}{1+e^{\theta(s)}}&\quad\textnormal{if}\leavevmode\nobreak\ \leavevmode\nobreak\ a=1,\\ \frac{1}{1+e^{\theta(s)}}&\quad\textnormal{if}\leavevmode\nobreak\ \leavevmode\nobreak\ a=2.\end{cases}$
@@
Note that there exists a one-to-one correspondence between soft-threshold policy parameterization and the hard-threshold function. That is, given a threshold parameter $\theta$, there exists a unique $f$ under the linear relation, $\theta(s)=f(y)-x,\leavevmode\nobreak\ s=(x,y)\in[L]\times[M]$.
In the parameterized setting the lagrangian value function for a given client $n$ with initial state distribution $\rho_{n}$, and $\lambda$ is given by,
(6) $J^{\pi_{\theta_{n}}}(\rho_{n};\lambda)=\mathbb{E}^{\pi_{\theta_{n}}}[\sum_{t=0}^{\infty}\gamma^{t}(c(s_{n}(t),s_{n}(t+1))+\lambda g(a_{n}(t)))]$
(7) $=\mathbb{E}^{\pi_{\theta_{n}}}[\sum_{t=0}^{\infty}\gamma^{t}c(s_{n}(t),s_{n}(t+1)]+\lambda\mathbb{E}^{\pi_{\theta_{n}}}[\sum_{t=0}^{\infty}\gamma^{t}g(a_{n}(t)))]$
Similarly, we denote the respective value function and the discounted constraint functions by
(9) $J_{c}^{\pi_{\theta_{n}}}(\rho_{n})\triangleq\mathbb{E}^{\pi_{\theta_{n}}}[\sum_{t=0}^{\infty}\gamma^{t}c(s_{n}(t),s_{n}(t+1)]$
(10) $J_{g}^{\pi_{\theta_{n}}}(\rho_{n})\triangleq\mathbb{E}^{\pi_{\theta_{n}}}[\sum_{t=0}^{\infty}\gamma^{t}g(a_{n}(t)))]$
It is instructive to view $J^{\pi_{\theta_{n}}}(\rho_{n};\lambda)$ as the value function with the cost function $\bar{c}(s_{n}(t),s_{n}(t+1))\triangleq c(s_{n}(t),s_{n}(t+1))+\lambda g(a_{n}(t))$. Therefore, we denote the corresponding advantage functions for $J^{\pi_{\theta_{n}}}(\rho_{n};\lambda),J_{c}^{\pi_{\theta_{n}}}(\rho_{n})$ and $J_{g}^{\pi_{\theta_{n}}}(\rho_{n})$ by $A_{\lambda}^{\pi_{\theta_{n}}}(s,a)$,$A_{c}^{\pi_{\theta_{n}}}(s,a)$, and $A_{g}^{\pi_{\theta_{n}}}(s,a)$ respectively. It is easy to show that for any state-action pair $(s,a)$, the following relation holds true, $A_{\lambda}^{\pi_{\theta_{n}}}(s,a)=A_{c}^{\pi_{\theta_{n}}}(s,a)+\lambda A_{g}^{\pi_{\theta_{n}}}(s,a)$.
The above parameterization readily enables us to use policy gradient methods. In particular, we use the framework of the Natural Policy Gradient (NPG) method for constrained markov decision processes (Ding et al., 2020a) to show finite time convergence results of our proposed algorithm.
Remarks: We would like to point out a few differences in the analysis of policy gradient methods whose convergence rates are summarized in (Agarwal et al., 2021a, Table 1). In particular, we restrict our attention to the class of soft-max parameterized policies owing to it’s similarity to the structure of the soft-threshold parameterization. Although, one can think of the soft-threshold parameterization as a special case, the convergence analysis does not carry over directly from it’s soft-max counterpart. Our reasons to choose NPG over the vanilla policy gradient method (PG) is multi-faceted. To begin with the finite convergence rate guarantees for the soft-max parameterization for the PG method requires one to optimize over a regularized optimization objective whose analysis use the special structure of the soft-max which is not true for the soft-threshold. This is due to the fact that the soft-threshold parameterization does not dependent on the action unlike the soft-max. The same is the case for the NPG-method where the Fischer-information matrix induced by the soft-max policy has a nice structure. However, with some additional algebra one can derive similarly simple parameter updates in the soft-threshold case for the NPG method. Moreover, the convergence guarantees of the regularized version of PG depends on a distribution mismatch coefficient with respect to the optimal policy which is unknown in general.
4.1. Convergence of NPG for Threshold Parameterization
Consider the following primal dual method to solve the minimization problem  1.

### v2

@@ 1. Introduction
Video streaming has grown in recent years to occupy about 70% of downstream mobile data traffic, with YouTube being the most popular (Sandvine, 2023). At the same time, increasing softwarization is taking place in both cellular and WiFi domains, allowing fine grained dynamic control of wireless resource usage to support mobile applications via intelligent control (Open Networking Foundation, 2021; Sunny et al., 2017). Indeed, intelligent control has been shown to enable dynamic priorities over the wireless edge (radio access network) for specific applications that require it (Ko et al., 2023).
@@
Figure 1. Feedback loop in a media streaming application. The states of the YouTube sessions and channel qualities are communicated to an intelligent controller that determines the service class for each session, accounting for resource constraints. The impact of this decision on the end-user QoE (reward) is communicated back to the controller.
@@ 1.1. Main Results
We first decompose the centralized CMDP problem across all clients into decentralized single-client problems using Lagrangian relaxation techniques, wherein a Lagrange multiplier acts as a price signal to enforce the constraint. As in several other queueing problem contexts, we find that the optimal policy of such a single-client problem has a threshold structure. Specifically, for any given stall count and constraint cost, the decision on whether to assign high priority service to a client only depends on whether or not its video buffer is below a fixed threshold.
We next consider whether we can provably learn the optimal policy. Here, the threshold structure of the optimal policy enables us to confine our search to learning the threshold parameter using a single neuron (soft threshold), as opposed to a complex function that needs to be approximated with a large neural network. We develop a primal-dual natural policy gradient algorithm, designed for learning such threshold policies under constraints. We show that the error in both the objective and constraints decay with time $T$ as $O(1/\sqrt{T}).$ Existing results on the convergence of policy gradient only apply to direct parametrization and soft-max parametrization. Hence, our methodological contribution is to prove the fast convergence of threshold-parametrized natural policy gradient, which applies to a variety of queueing systems.
We next develop a simulation environment to train optimal policies. We show that the single-client decomposition approach learns at about four times faster than centralized learning while attaining a similar performance. Furthermore, the single neuron structure implies that the inference time for decision making is only about 10 to 15 $\mu$s, as opposed to 50 to 60 $\mu$s taken by an unstructured policy, i.e., the approach is much lower in computational requirements and can be run in a realtime manner. We also develop a heuristic index-variant that has the advantage of being applicable in spite of a changing number of clients or variability in channel quality.
@@ 1.2. Related Work
There is an extensive amount of work on CMDPs focusing on the problems of control (Altman, 1999). In the context of media streaming over wireless channels, several works identify structural properties of the optimal policy (ParandehGheibi et al., 2011; Singh and Kumar, 2019; Hsieh and Hou, 2018) that often take a threshold form. While (ParandehGheibi et al., 2011; Singh and Kumar, 2019) consider maximizing QoE for a finite system, and (Singh and Kumar, 2019) utilizes a dual-based decentralization approach, (Hsieh and Hou, 2018) focuses on the heavy traffic limit.
@@
Reinforcement learning approaches to find the optimal policies for CMDP problems are well studied in recent literature. In the context of a tabular setting, recent works focus on characterizing the regret (Liu et al., 2021; Wei et al., 2022; Bura et al., 2022) or the sample complexity (HasanzadeZonuzy et al., 2021a, b; Vaswani et al., 2022) of learning algorithms. There also a number of works on policy gradient approaches for CMDP (Achiam et al., 2017; Zhang et al., 2020; Ding et al., 2020). However, these works are for the general CMDPs and they do not exploit the structure of the problem to learn a threshold policy with global convergence guarantees.
Other work that considers a learning approach to streaming is (Roy et al., 2021), which develops a two time-scale stochastic approximation algorithm (Borkar, 2009) for an unconstrained MDP, and shows certain structural properties of the value function (unimodality), which suggests that a gradient algorithm might converge to the optimal. However, optimality of the fixed point of gradient descent is not shown. We take a very different approach via constrained natural policy gradient, and obtain guarantees for fast convergence to the optimal policy.
@@
A variety of systems have been proposed to improve the QoE performance of video streaming. In software defined networks (SDN), assigning the video streaming flows to network links according to path selection algorithms is considered in (Jarschel et al., 2013). VQOA (Ramakrishnan et al., 2015) and QFF (Georgopoulos et al., 2013) employ SDN to monitor the traffic and adapt the bandwidth assignment of video flows to achieve better streaming performance in the home network. Reinforcement learning has also been applied to a variety of video streaming applications (Mao et al., 2017; Bhattacharyya et al., 2019), which show significant improvements over existing methods. (Mao et al., 2017) proposes an algorithm Penseive, to train adaptive bit rate (ABR) algorithms to optimize the QoE of the clients.
Finally, (Bhattacharyya et al., 2019) considers the problem of service prioritization for streaming, and shows empirically that off-the-shelf model-based and model-free RL approaches can result in major improvements in QoE. While completely heuristic in its approach, the work motivates us to consider the problem of optimal policy design, while exploiting problem structure. Thus, we provide a simple, structured RL algorithm with provable convergence guarantees, which is empirically able to attain the same performance as more complex deep learning approaches, while only using significantly fewer data samples, and with much less compute needed during runtime.
@@ 2. PROBLEM FORMULATION
We consider a discrete time system with $N$ clients that are connected to the AP and compete for resources. At any time $t$, a media server sends some number of media packets to the AP, which in turns forwards them to the appropriate client, which buffers up this content, while continually attempting to play out a smooth stream. The media playout is said to stall when the media buffer is empty, and is an undesirable event. Stall events cause end-user dissatisfaction, and we will utilize a standard model of this disutility in the video streaming context.
@@
We assume that $\mu_{n,1}>\mu_{n,2}$, for each $n$. This means that when the client is assigned to high priority service, its service rate is higher than that of the low priority service. The joint system action space is $\mathcal{A}\triangleq\otimes_{n=1}^{N}\mathcal{A}_{n}$. Note that we could easily include video bitrate adaptation into the model by simply increasing the action space to multiple dimensions. This causes no changes to the approach. We now describe the dynamics of the system.
@@
State Transition Structure: The system state has two component-wise transition processes. The first process corresponds to the media buffer evolution, while the second corresponds to the stall count evolution. We assume that the media playback process of client $n$, denoted by $B_{n}(t)$ is distributed according to a Bernoulli distribution with mean $\beta_{n}<\mu_{n,1},$ i.e., stall-free playout is possible if the client is given a high priority service. Furthermore, a stall event corresponds to a transition from a non-zero media buffer state to a zero media buffer state (the buffer cannot be negative).
We assume that the media stream that is currently being played by the client may be terminated at any time with probability $\alpha>0$ by the end-user (i.e., the user simply stops the playout), which causes a reset of the client state to $(0,0).$ Hence, the possible state transitions from $s_{n}(t)$ to $s_{n}(t+1)$ are
@@
where $x^{+}=\max(x,0)$, $y_{n}^{\prime}(t)=y_{n}(t)+\mathbbm{1}\{x_{n}(t)=1\}$, $P_{n,1}(a)=1-P_{n,2}(a)-P_{n,3}(a)$, $P_{n,2}(a)=\mu_{n}(a)(1-\beta_{n})$, $P_{n,3}(a)=(1-\mu_{n}(a))\beta_{n}$, for $a\in\{1,2\}$, and for every $n$. The state transition probability of the $n$-th client is denoted by $p_{n}(s^{\prime}|s,a)$. The starting state distribution of the $n$-th client is denoted by $\rho_{n}$. Note that we can include client-specific transition probabilities based on their individual channels without changing the transition structure.
@@
Note that we choose to use a cost function, rather than a QoE-based reward function for consistency with a generic MDP/RL approach that focuses on cost minimization. It is easy to translate the instantaneous cost to an instantaneous reward by simply multiplying by $-1$ and maximizing. As pointed out above, we can easily include the video bitrate (corresponding to the video resolution) in this cost function but choose not to do so for simplicity.
@@
$\min_{\pi\in\Pi}\mathbb{E}^{\pi}\left[\sum_{n=1}^{N}\sum_{t=0}^{\infty}\gamma^{t}c(s_{n}(t),s_{n}(t+1))\right]$
@@
(1) $&\min_{\pi\in\Pi}\mathbb{E}^{\pi}\left[\sum_{n=1}^{N}\sum_{t=0}^{\infty}\gamma^{t}c(s_{n}(t),s_{n}(t+1))\right]\\ &\text{s.t}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{E}^{\pi}\left[\sum_{n=1}^{N}\sum_{t=0}^{\infty}\gamma^{t}g(a_{n}(t))\right]\leq\bar{K},\end{aligned}$
@@
An immediate consequence of Assumption (2) is that there exists a primal-dual pair $(\pi^{*},\lambda^{*})$ achieving strong duality. For the ease of notation, we denote the individual value and discounted constraint functions of client $n$ under policy $\pi$ by,
(2) $J_{c}^{\pi}(\rho_{n})$ $\triangleq\mathbb{E}_{s_{n}(0)\sim\rho_{n}}\left[\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{n}(t),s_{n}(t+1))|s_{n}(0)\right]\right],$
(3) $J_{g}^{\pi}(\rho_{n})$ $\triangleq\mathbb{E}_{s_{n}(0)\sim\rho_{n}}\left[\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}g(a_{n}(t))|s_{n}(0)\right]\right],$
@@
Given that assumption (2) holds true, then, $D(\lambda^{*})=\sum_{n=1}^{N}J_{c}^{\pi^{*}}(\rho_{n})$
However, finding the optimal joint policy even for a fixed $\lambda$ in the dual function can be computationally hard because the complexity increases exponentially with the number of users. It can be shown that the joint problem can be decomposed in to $N$ smaller problems (Singh and Kumar, 2019).
@@
The proof is similar to that of (Singh and Kumar, 2019, Theorem $1$), for completeness, we provide the details in Appendix A.2.
Note that Theorem (1) states that it is sufficient to look at decentralized class of policies of the form $\pi=\otimes\pi_{n}$.
3. Structure of the optimal policy
@@ 3.1. Threshold Structure
We will use the following results to prove the above theorem. Since the transition function of all the clients have the same properties (even though they may not be identical), we drop the the client index $n$ for ease of notation, and denote the corresponding optimal value function for state $s$, and parameter $\lambda$ by $J^{*}(s;\lambda)$.
@@
The proofs of Lemma 2 and Lemma 3 are given in Appendix A.3. We now present the proof of Theorem 2.
@@ 4. RL for Learning the Optimal Threshold
In this section, we develop a primal-dual algorithm to solve the constrained Markov decision process problem, and provide results for its convergence to the globally optimal policy. Our implementation of the proposed algorithm follows the methodology of actor-critic algorithms with function approximation. To begin with, we slightly modify our hard-threshold policy to a parameterized soft-threshold policy. For the soft-threshold parametrization, the probability of taking action $a$ in state $s$ corresponding to a threshold parameter vector $\theta$, $\pi_{\theta}(s,a)$ is defined as
$\pi_{\theta}(s,a)$ =$=\begin{cases}1-\frac{1}{1+e^{\theta(s)}}&\quad\textnormal{if}\leavevmode\nobreak\ \leavevmode\nobreak\ a=1,\\ \frac{1}{1+e^{\theta(s)}}&\quad\textnormal{if}\leavevmode\nobreak\ \leavevmode\nobreak\ a=2,\end{cases}$
@@
Note that there exists a one-to-one correspondence between soft-threshold policy parametrization and the hard-threshold function. That is, given a threshold parameter $\theta$, there exists a unique $f$ under the linear relation, $\theta(s)=f(y)-x,\leavevmode\nobreak\ s=(x,y)\in[L]\times[M]$.
In the parameterized setting, we denote the respective value function and the discounted constraint functions by
(6) $J_{c}^{\pi_{\theta_{n}}}(\rho_{n})\triangleq\mathbb{E}_{s_{n}(0)\sim\rho_{n}}\left[\mathbb{E}^{\pi_{\theta_{n}}}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{n}(t),s_{n}(t+1)|s_{n}(0)\right]\right]$
(7) $J_{g}^{\pi_{\theta_{n}}}(\rho_{n})\triangleq\mathbb{E}_{s_{n}(0)\sim\rho_{n}}\left[\mathbb{E}^{\pi_{\theta_{n}}}\left[\sum_{t=0}^{\infty}\gamma^{t}g(a_{n}(t)))|s_{n}(0)\right]\right]$
Similarly, the Lagrangian value function for a given client $n$ with initial state distribution $\rho_{n}$, and $\lambda$ is given by
(8) $J^{\pi_{\theta_{n}}}(\rho_{n};\lambda)\triangleq J_{c}^{\pi_{\theta_{n}}}(\rho_{n})+\lambda\left(J_{g}^{\pi_{\theta_{n}}}(\rho_{n})-\frac{\bar{K}}{N}\right).$
It is instructive to view $J^{\pi_{\theta_{n}}}(\rho_{n};\lambda)$ as the value function with the cost function $\bar{c}(s_{n}(t),s_{n}(t+1))\triangleq c(s_{n}(t),s_{n}(t+1))+\lambda(g(a_{n}(t))-K/N)$. Therefore, we denote the corresponding advantage functions for $J^{\pi_{\theta_{n}}}(\rho_{n};\lambda),J_{c}^{\pi_{\theta_{n}}}(\rho_{n})$ and $J_{g}^{\pi_{\theta_{n}}}(\rho_{n})$ by $A_{\lambda}^{\pi_{\theta_{n}}}(s,a)$,$A_{c}^{\pi_{\theta_{n}}}(s,a)$, and $A_{g}^{\pi_{\theta_{n}}}(s,a)$ respectively. It is easy to show that for any state-action pair $(s,a)$, the following relation holds true: $A_{\lambda}^{\pi_{\theta_{n}}}(s,a)=A_{c}^{\pi_{\theta_{n}}}(s,a)+\lambda A_{g}^{\pi_{\theta_{n}}}(s,a)$.
The above parametrization enables us to use policy gradient methods. In particular, our approach to prove convergence is inspired by the framework of the Natural Policy Gradient (NPG) method for Constrained Markov Decision Processes (Ding et al., 2020).
4.1. Convergence of NPG for Threshold Parametrization
Consider the following primal-dual method to solve the minimization problem (1).